Verify the MIRR is 9.29% given cash flows in years 1 and 2 of $1,000 each, cost of asset of $1,800, and cost of capital of 15%.
Answer #1

MIRR = (FV of Cash Inflow/PV of Cash outflow)1/n - 1

FV of Cash Inflow = 1,000(1.15) + 1,000 = $2,150

MIRR = (2,150/1,800)1/2 - 1

MIRR = 9.29%